OpenStudy (anonymous):

Can someone help me identify the vertex and x-intercepts of f(x)=-(x^2+x-30) and then check the results by writing the quadratic function in standard form??

OpenStudy (phi):

there are a few ways to do this: 1) factor the quadratic to find the x-intercepts. the x value of the vertex will be exactly in between them. then use the equation to find the y value of the vertex

OpenStudy (phi):

2.) complete the square, so that you can write the equation in "vertex form"

OpenStudy (anonymous):

which way do you think would be better?

OpenStudy (phi):

3) write the equation in standard form (i.e. distribute the -1) and then use x value of vertex = -b/(2a) where a,b (and c) are the coefficients y= ax^2 + bx+c

OpenStudy (phi):

I would try to factor the quadratic. list the factors of 30

OpenStudy (anonymous):

I got (1,2,3,5,6,10,15,30)

OpenStudy (phi):

1,30 2,15 3,10 5,6 are all of them the minus sign on -30 (the last number) means the factors have different signs. the plus sign on +1x means the larger one is positive. you want the pair that gives you +1 when you add the pair (with the larger of the pair +, and the smaller of the pair -)

OpenStudy (anonymous):

could it be +6 and -5?

OpenStudy (phi):

yes, that means x^2+x-30 factors into (x-5)(x+6) remember we have that minus sign out front: f(x)= -(x^2+x-30) we get f(x)= -(x-5)(x+6)

OpenStudy (phi):

the x-intercepts make f(x) zero. in other words, solve for x in x-5=0 and x+6= 0

OpenStudy (anonymous):

x=5 x=-6 but what about the negative sign out front?

OpenStudy (phi):

the minus sign out front "flips" the parabola so it looks like a frown... but the x-intercepts are the same.

OpenStudy (anonymous):

oh okay, so the x intercepts would still be 5 and -6?

OpenStudy (phi):

yes. Here is a graph with and without the minus sign the blue one is yours

OpenStudy (anonymous):

what formula would i use to find the vertex?

OpenStudy (phi):

the x value of the vertex is half-way between the x values of the x-intercepts

OpenStudy (anonymous):

so its a number in between 5 and -6?

OpenStudy (phi):

the number *exactly* in the middle between 5 and -6

OpenStudy (phi):

one way to find the middle number is find the average of 5 and -6

OpenStudy (anonymous):

-0.5?

OpenStudy (phi):

yes, that is the x value of the vertex. to find the y value, use the equation f(x)=-(x^2+x-30)

OpenStudy (anonymous):

I got 29.25.. is that correct?

OpenStudy (phi):

your calculator is broken -( (-½)^2 + -½ -30) -( ¼ - ½ -30) -(-¼ -30) -(-30.25) 30.25

OpenStudy (anonymous):

oh i used .5 instead of -.5!

OpenStudy (phi):

the vertex is at (-0.5,30.25) the x-intercepts are at x= -6 and x=5

OpenStudy (phi):

for this part quadratic function in standard form?? I think we distribute the minus sign to get f(x) = -x^2 -x +30

OpenStudy (phi):

I don't know how that "checks" anything, unless they wanted you to write the equation in vertex form y= a(x-h)^2 + k where (h,k) is the vertex and then change it into standard form and show you get the same equation btw, vertex form is y= -(x - (-0.5))^2 + 30.25 y= -(x+0.5)^2 + 30.25

OpenStudy (anonymous):

thank you so much, i really appreciate it!! :)